Home
FM ILT - PACS at Leuven
Contents
1. Limit switch investigation this seems to be a hardware failure indeed 1 e either the line to the switch or the switch itself 1s broken the redundant switch seems to be ok For the further tests until Christmas we will just 1gnore this switch Cryostat started pumping on Saturday very cold now need to adjust to 1 6K Liquid Nitrogen has to be refilled will be finished by 10 30 today First cooler recycling is planned for tomorrow PCSS This week MW will work on detector selection table issues Other Over the weekend there was a telemetry loss event PACS was sent into save mode by the event monitor HF and ErW to analyse Today FOV scans move the chopper in small steps thermal behaviour tests use only 5Hz chopping so the current chopper settings should be ok for this week If confirmed we will do over night tests to further optimise the chopper settings In order to concentrate fully on the bolometer tests where we need the French connection which can only stay until Wednesday we will shift the thermal behaviour test to later this week analyse the telemetry loss event HF ErW adjust cryostat and refill nitrogen GJ get input from L Barl for Hell Ge Ga settings do the SFT Hell 12 000 11 000 4 ood ood ood ooo ooo ood 2 Odo 4 ooo 0 0 0 0 4 52 Hi BM Dog UD OO od uw oc ooo 4 60 ood ooo ood ooo O00 0 9 000 4 10 000 11 000
2. 12 000 13 000 chepper r adsute of file ne 10004 oe Cy Om lk e HB o2500 aana a8an u aana amp as un sd 5500 a amp nan SeS500 index B chopper readouts of file no 52 4 1 deg chopping x axis is in ms 11900 F 11500 12000 4 11800 12100 11700 12200 4 11600 ia ia 11500 12300 o 9 B 11400 B 12400 E a AH ri d 11300 d 12500 3 11200 12600 4 n n H H E 11100 8 12700 J i 11000 8 r i 12800 4 10900 a 12900 E 10800 13000 J 10700 YEAH 13100 J 10500 713200 7 10400 13300 t t t t t t t 86200 86300 86400 86500 86600 86700 86800 86900 87000 87100 87200 87300 87400 91500 91600 91700 91800 91500 92000 92100 92200 index index chopper readouts of file no 52 chopper readouts of file no 52 4 1 deg chopping zoom x axis is in ms 4 1 deg chopping zoom x axis is in ms 56 chopper readouts of file no 56 chopper r deuts of file no 20 000 17 500 15 000 4 12 500 10 000 7 500 4 5 000 4 2 500 o 2 500 4 5 000 7 500 4 10 000 4 12 500 15 000 17 500 20 000 4 22 500 4 81000 4 1 t t T t t t 81500 82000 82500 83000 83500 84000 84500 85000 index B chopper readouts of file no 56 6 1 deg chopping x axis is in ms dl 21550 1 22800 22850 21500 7 22900 21450 4 2295
3. in Volt units which then in the system get converted into raw values Small biases were converted into zero and the command rejected For the time being we can circumvent this by using the bias tables and commanding raw units directly ErW to analyse MIB or Test Control problem Today investigate switch off involve Stefano after switch on prepare phot mode do the photometer thermal behaviour tests in parallel investigate the problem DC reported and when this is solved the push button installation 20061107_10 BLUE HEATER TEST SSS ae ee a Sm Se mmm EE EUM E mE mm m m m mmm aea me a me rm im mm mL RI ee m mm m mm m UE Nm fg eer EM m S m mmUmS mmm RARE ESAE qmm Rm EE NU mmm gm M SO RU mm m ES m RES SUR mm E mS NON m M BIN mU S m E E Emm mS MORE m e M E mmm E P E MS ES m mm E Em mmm LEE T b d 2 2 2 2 2 2 2 b 2 2 2 222 2 2 2 2 d2 2 2 2 2 22 2 2 2 2 2 b 2 2 2 2 2 22 22 2d2 2 222 2 2 2 2 2 2 2 Lbe Jd E EL H 2 T 2 Lb 22 2 d2 22 2 2 2 2 2222 2 2 bh 2 2 2 2 22 22 2d2 2 2 2 2 2 2 ER Bb 2 uu ee ee ee ee nn gp o m a a embannan eee ee eee eee ee ee eee bennnk 0 0 0 5 0 6 0 9 I2 1 5 1 8 2 1 DM DECB HEAT C mA PR reported yesterday email on the heater settings the setting to send the blue detectors to 2 5 K is 1 5 mA i e 310 in raw commanding units A CUS modul
4. overflow problem We need to assign people to check these kind of statuses and run the FM ILT routine checks as explained on the wiki pages RV and B Alt might be willing to take over this task for a while We then moved on to the cooler recycling test which was originally planned for today The pre conditions were ok except for the evaporator temperature which was at 2 08K rather than the required 2K but this has always been the case in the past Duration 2 2hrs 2hrs 12min The evaporator temperature did not drop fast enough and at the end was just above the required 400mK which then triggered a save mode The cooler recycling has to be modified extend the time at the end by 15 minutes or so and we need to pump down as much as possible We need to investigate this further probably there should be no problem in the Herschel cryostat After recycling we planned to prepare photometry mode when there was another TM loss event 18 30 This one caused a complete controlled switch off of PACS autonomy function Before we switch on again we will investigate what happened was it related to a specific possibly wrong command etc PCSS In the morning ErW tried to install the push button for CUS scripts but Test Control did not work on PACS 5 ErW is investigating the problem For the time being we have to use the Test Control server on PACS i e just like before DC reports a problem he tried to command bias values
5. that the lag may be non linear tracking this down will require much more work lag the shift in frames needed to relate the signal between a reference pixel and the pixel whose data 1s being analyzed Applying this lag means that both pixels were looking at the same part of the FOV and hence the same signal The map is built up using this information We noted that the finetime in the data had quite a lot of discontinuities While this did not affect the analysis this is definitely worth investigating further While at first look the pixel to pixel signal offsets looked better for the FM than COM there is still very significant differences between pixels Our first attempt at ignoring these offset for a quick and dirty map produced horrible final images Further work derive the chopper calibration using the edge of the calibration blackbody investigate the lag Incorrect lag values wash the data out Even for quick analysis we need the best value we can derive redo the maps using the new calibration redo the FOV map using CM or another mapping algorithm cheers Babar PACS FM ILT Daily morning briefing Date 17 November 2006 8 30 Friday Running Number 013 Participants ICC MPE GJ Chr H Absent AlPog Minutes ES Status Cooler Cooler hold time was 48 hours below 300mK Bolometer tests were continued all planned tests completed This involved the OGSE BBs and OG
6. 0 4 23000 21400 F 23050 4 21350 4 23100 7 2 231504 ARTT A 23200 4 d Ww 21250 23250 9 23300 poe E 23350 21150 4 23400 4 H 23450 21100 4 B 23500 21050 4 E 23550 4 23600 4 21000 23650 4 20950 4 23700 23750 4 20900 23800 4 i t t t t t t 77000 77200 77300 77400 77500 77600 77700 84000 4100 4200 4300 4400 84500 4600 4700 index index m chopper readouts of file no 56 W chopper readouts of file no 56 8 1 deg chopping zoom x axis is in ms 6 1 deg chopping zoom x axis is in ms PACS FM ILT Daily morning briefing Date 14 November 2006 9 15 Tuesday Running Number 010 Participants ICC 9 MPE GJ AlPog B Alt RV MS KO O Boulade N Billot Absent HF Minutes ES Status SFT Hell was executed successfully Analysts have been informed and are working on the data analysis During the sorption cooler functional check ou the stabilisation time of the cold electronics at 2K takes longer than at 4K When using the same biases at 2K and 4K there is not enough current in the electronics Towards the end some science data were lost Possible reason an interference with the data base replication TBC For the time being data base replication will only be done during spare times of the ILT when no commands are running Compression was not as efficient as required The TM problem over the weekend was a disk
7. CS Trend Analysis Plot No 2 Source home tmueller FILT ThermalBehaviour Phot 20061114 02 un a E eng K B b 4 4 L TEMP S U ST eng X BDL L TEMP F2 Hd 4 a N 1 34 1 t 1 542 210 000 000 000 1 542 212 500 000 000 Time microseconds eet eng K 8D Time BOL TEMP EV SWT eng K B Time BOL TEMP FPU ST eng K B Time BOL TEMP SP eng K Time BOL TEMP TS eng K PACS Trend Analysis Plot No 1 Source home tmueller FILT ThermalBehaviour Phot 20061114 02 tm en 0 2975 4 0 2950 4 0 2925 L TEMP FPUZ 8D 0 2900 4 K 0 2875 eng 0 2350 4 a 2825 a 2200 4 L_T amp Me_FPul 8D 0 2775 4 K 0 2750 eng 0 2725 4 2700 i n 1 542 210 000 000 000 1 542 212 500 000 000 Time microseconds BDL TEMP amp V a Bi Time BOL TEMP EV eng K B Time BOL TEMP FPUI eng K B Time BOL TEMP FPU2 eng K PACS FM ILT Daily morning briefing Date 16 November 2006 8 30 Thursday Running Number 012 Participants ICC MPE GJ AlPog Chr H W Sch Absent ES Minutes HF ES B Ali Status Cooler is still recycled science data is still flowing with detectors permanently biased Only one 10 15 min interruption Several reference measurements done in all sequence modes to obtain clean data for SPU processing examples Problem with bias tables has been solved no
8. PACS FM ILT Daily morning briefing Date 13 November 2006 8 30 Monday Running Number 009 Participants ICC 9 MPE GJ AlPog B Alt RV MS KO O Boulade N Billot Absent OHB part time Minutes ES Status Grating was tested with 5 and 6 deg s data to be analysed before we try 8 deg s Analysis script was updated to avoid Jython loops Diagnostic HK was analysed and made available to Benoit Until we get confirmation that we can move faster we will use 4 deg s and 1deg s for homing CSs all data taken MPIA to analyse TM wants to extend the thermal behaviour tests a bit to assure that the full ballistic heating phase is covered Chopper MPIA left two parameter sets for phi lt 4 1 deg and phi 8 1 deg and the angular calibration report including the maximum allowed positions These parameter sets are not yet optimal long transition times asymmetries two sets for different ranges but show no over undershooting and swing in anymore and have low noise The settings should be good enough for the bolometer tests that are planned for this week see below HF suggests and will send an email to MPIA for discussion a suite of over night tests trying out ten thousands of controller parameter sets for further optimisation A Mazy will look into options to modify the s w to account for the asymmetries and to allow different chopper parameter sets in the commanding There will be a telecon later this week
9. SE chopper The OGSE chopper reaches much faster chopping frequencies now 2 3Hz Detector selection table tests ok just some confusion because the new tool by Roland O seems to be a detector de selection tool Note all existing detector selection tables are listed on the Wiki pages OBCP tests all existing photometer OBCPS were tested with realistic parameter settings Everything worked fine except for the freeze frame mode This probably needs a different chopper parameter set commanded too slow TBC Buffer transmission mode photometry tests ok The autonomy function for the bolometer temperatures has a s w problem it works when the temperatures are OOL right away but it does not react when they develop slowly into OOL during an observation This 1s a serious problem for overnight tests but Stefano is informed and is working on it For the time being we must avoid over night tests until this 1s fixed SFT chopper repeated Some manual commanding of the spectrometer detectors with the current capacitor setting 1 e more noise and the current SPU s w version the SPU goes into auto stop 16 sample long sub ramps are too heavy for the CPU load DPU s w upload MIB update Cryostat LO still low but level meters show zero These meters are not very sensitive to low gas flow GJ installed a new gas meter for very low flows We will get warmer soon PCSS Automatic procedure to check data base size instal
10. able to look at one file and one filter data From our estimation a proper quick report will likely take another couple of days to finish This is much slower progress than I anticipated however see below The enclosed tarball has the figures that might be of interest along with an file explaining what they are I ll repeat the contents of the file below We had computation troubles with a 8Gig ram 8 CPU 64 bit machine Decompression was easy Analysis particularly cube analysis brought the machine to its knees In the end we had to increase the swap memory to 32 gig in order to do the analysis This may be due to configurations on our end and we will certainly explore that We believe that the chopper throw on sky vs CPR value may be incorrect or atleast inconsistent Tracking this down and actually believing it took a little while but this explains why our first few FOV maps looked much to smoothed out with the calibration sources This is a crucial first step that we had to derive in order to do the shift and add mosaicking This was probably the biggest hurdle See the enclosed file for more details Another time consuming matter was finding the lag needed to relate the signal from one pixel to another in a given row Our empirical estimates suggested values between 14 and 17 readouts Theoretically it should be 10 readouts However the final maps did not look perfect with either one of these lags We suspect
11. e named PACS Spec Heat Setup OBS exists already in order to do that 20061106 Blue Heater Characterisation 10 S00 1800 2 00 3600 4500 5400 6300 7 200 8100 Time sec plot of the temperature vs time PACS FM ILT Daily morning briefing Date 15 November 2006 8 30 Wednesday Running Number 011 Participants ICC 2 MPE GJ AlPog B Alt RV MS KO O Boulade N Billot Chr H W Sch Absent Minutes ES Status TM loss investigation the event was caused by the DPU which has a known OBSW bug This will be fixed with the next DPU s w upload tomorrow Cooler recycling repeated GJ improved the cool down procedure temperatures looked better this time including T evap which is still not quite at the temperature that was measured at Saclay The required 400mK were reached just one minute before end of the extended script We are now 20 hours into cooler recycling and stll cold Set up photometry with the new set of schopper parameters worked fine Bias fine tuning a slightly different bias table than tin the user manual is now used AI MS will update the user manual asap Detector sorting a problem of the past 1 e 1s ok Another TM los occurred probably due to a timing problem Will be corrected today The cooler did not suffer during the time when PACS was switched off Thermal behaviour photometry TM showed three figures see below Filter Wheels and Calibration Sources i
12. etector selection table OGSE chopper together with bolometer signal two OGSE BBs in use OBCP tests Tutti Frutti with all chopper positions adapted to FM DPU OBSW upload and MIB upgrade Thermal behaviour Spectroscopy SFT chopper Hell Report from over night analysis of FOV scan at NHSC FM chopper calibration used for FOV maps 40 24 32 16 Angle degrees 8 0 24 16 i 2 000 18000 9000 0 9000 18000 27 000 Aua Readout EPR 2000 2000 0 Signal Pixel 12 32 6300 5600 4900 4200 3500 2800 210 0 40 80 120 160 200 240 280 320 6 4 2 0 2 4 6 Col No Chopper throw arcmin Layer 0 m N mM N re AX A i Un e e e N l e e e T 2000 4000 0 Signal Pixel 12 32 2000 0 2000 4000 4000 6 4 2 0 2 4 6 28000 21000 14000 7000 Chopper throw arcmin column 0 column 63 O 7000 14000 21000 CPR column 63 column 0 Figure explanation fm cpr vs ang jpg This figure shows our chopper calibration This is essential for shifting and coadding data The blue cyan and green curves show the calibration derived by MPIA from laser data PICC MA TR 009 for the three different ranges of chopper angles The final curve we used employs combining the three segments into one The boundaries are marked with the dashed line Reading from left to right you should look at the cyan line switch to blue curve at the
13. led The procedure for the merging of the x y stage housekeeping and the science data has been developed and tested There is a time mismatch which shall be solved today The high SPR on frames generation will be solved by today evening too A demo of data base tools is ready can be given at any time where there is a short break in the tests Monday Other From the flux ratios we see HF gets the suspicion that the photometry filters of PACS are swapped i e A long wave filter B short wave filter To be followed up Today refill and pump 4 hours thermal behaviour test spectroscopy first light in spectroscopy
14. negative boundary shown in dashed red switch to green curve at the positive boundary shown in dashed gold sigbs chopthrow row12 eps This figure shows the signal for blue short for a central pixel 32 plotted vs the chopper throw angles in arcminutes on the sky The CPR readouts are converted to chopper angle on the sky using the calibration from the figure above The magnification used 1s 3600 80 69 see PICC MA TR 009 Notes The available FOV implied by this figure is 6 arc minutes However see our notes below for the chopper calibration sigbs firstlastcol chopthrow row12 eps This figure shows simulatneous plots of the first and last column First column is shown in blue Last column is shown in red The chopper throw is calculated from CPR as for the figure above Notes This is our first indication that the chopper calibration may not be accurate Note that the available FOV the flat section where there 1s no light from the calibration sources is wider by 1 arc minute for the first column than for the last column The first column covers the negative throws the last column covers the positive throws While no symmetry is expected the flat section should have the exact same width for both implied FOVs see also below The simplest explanation is that either the negative or the positive chopper angle to sky throw are inconsistent This figure also suggests that the right calibration source CS2
15. ntroduce a lot of heat as expected But the bolometer temperatures are not affected a lot There are some strange spikes on the 300 mK level especially at the end and there is some risk that an autonomy function could be triggered by this Has never been seen before Electrical noise The instrument was left on over night in high gain Cryostat Helium arrived yesterday with some delay because the truck broke down on the Autobahn Supply seems to be secure for the next two weeks Today Fix the bias tables in Volts problem Postpone the upload of the 2 new MIBs OGSE and PACS to tomorrow FOV scans OGSE characterisation including x y stage At 17 00 telecon with NHSC to hand over the FOV scan data analysis PACS Trend Analysis Plot No 3 Source home tmueller FILT ThermalBehaviour Phot 20061114 02 tm eee a 5 104 05 4 sw PHDT T amp E uU 5 00 4 DM r I L 95 4 90 K 4 357 eng 4 80 4 75 e 70 4 657 e 60 t DM FP U T2 TEMP 55 4 e K e 50 4 45 F eng re a4 a0 1 35 1 542 210 000 000 000 1 542 212 500 000 000 Time microseconds Wi Time DM CAL SRC TEMP eng K Bl Time DM_CHOPPER_TEMP eng K B Time DM_FPU_T1_TEMP eng K Time DM_FPU_T2_TEMP eng K BI Time DM FW PHOT TEMP eng K M Time DM FW SPEC TEMP eng K E Time DM GRATING TEMP eng K DM s U T1 TEMP PA
16. starts to hit the right edge of the array at arc minute This clearly cannot be the case sigbs firstlastcol cpr row12 eps This figure is the same as sigbs firstlastcol chopthrow row12 eps except that instead of arc minutes on the sky the x axis is now CPR Notes The width of the flat part no signal 1s now the same in terms of delta CPR values that it spans Ie The range of CPR values for the flat part 1s consistent The fact that this 1s inconsistent when the same signal 1s shown vs sky angle indicates to us that the calibration 1s inconsistent NOT SHOWN The fov obtained by using the chopper calibration discussed above We noticed a significant smearing of the signal and will work to improve on this by improving our estimate of the chopper calibration fov pg This is the final FOV map obtained by using a constant value of 199 CPR values pixel This approach towards making a FOV 1s clearly not accurate but yielded much better results than shift and add approach using the chopper calibration above fov row12 eps A row cut through fov jpg Dear All We had a long day of analysis and I m almost too tired to write this e mail The data had a lot more surprises in it than we thought and I will try to bullet ize the main findings points from our session today Despite what you read below it was very exciting to work on the first light data Too bad we missed the Whisky n the end we were only
17. w also engineering tables can be used FOV scans have been executed for dark dark CSs OGSW 22K both cryostat windows from to both extreme chopper positions in both filters At this point we had a little first light celebration Cui Look Analyzis lt 3 jmi heer me ee eee dam ie lle ee eee Stes No Plczdmamarun iP poe Slade degzicibce F Phoiererts lags isisi megs Sole gaib drin barm ane Both flip mirrors moved correctly in the OGSE Bolometer behaviour has been recorded also during transition stabilization phases BB turn on change of windows etc basic manual x y stage commanding but no raster yet Zero position for x y stage is not yet defined Al on GJ to provide the measured zero position for FM ILT to ErW Small problems of the MIB written input required since it also affects DPU OBSW Generic CUS commanding exercised all automatic transitions and also more manual interactive scripts Not all data is at Leuven due to interferences of copy process with real time activities on PACS1 AI on ErW to take care that all data goes to Leuven FOV Row 12 Status Cryostat n general o k LO level is running on the low side now could happen that we have a problem tomorrow and a refill 1s required Whatever next week shall start with a cooler recycling in the morning Status PCSS nothing in particular Plans for Today Continued photometer tests buffer transmission mode d
Download Pdf Manuals
Related Search
Related Contents
Manual do usuário - Metto Ferramentas TR-20 User Manual - NRCS Irrigation ToolBox 操作方法 Samsung CB-21N30MJ User Manual Emerson Fisher 3582 Data Sheet KVM-0223 Cable KVM Switch, 2 Ports, VGA, USB, Audio User Manual PCグリーンラベル環境性能の可視化に関する規程 Ultron US-21 TecnoaXess© – Manuale Utente betriebsanleitung tuner evolution t2 rds Copyright © All rights reserved.
Failed to retrieve file